# Symmetry transformation group arising from the Laplace–Runge–Lenz vector

### v2

@@ 5. LRL symmetry transformation group
This symmetry transformation group acts as isometries in the space of constants of motion $(\vec{L},\vec{M})$, where $\vec{L}\cdot\vec{M}=0$ and $|\vec{L}|^{2}-{\rm sgn}(E)|\vec{M}|^{2}=\kappa^{2}/(2|E|)$, with $|\vec{L}|\neq 0$ and $|\vec{M}|\neq 0$. When $E<0$, the space is isomorphic to $S^{2}\times S^{1}\times S^{1}$, which has the symmetry group $SO(4)$. The space is isomorphic to $S^{2}\times S^{1}\times\mathbb{R}$ when $E>0$, and the symmetry group is then $SO(3,1)$.
@@
The LRL dynamical symmetry transformations (5.8) have the physical meaning that they preserve energy, while angular momentum and the LRL vector are changed in a specific way that depends on the energy. In the case $E<0$, the vectors $\vec{L}$ and $\vec{M}=\sqrt{2|E|}^{-1}\vec{A}$ are rotated, preserving $|\vec{L}|^{2}+|\vec{M}|^{2}$ and $\vec{L}\cdot\vec{M}=0$, along with the period $T=2\pi/\sqrt{2|E|}$ and semi-major axis $a=\kappa/(2|E|)$ of the elliptical orbit. Likewise in the case $E>0$, $|\vec{L}|^{2}-\frac{1}{2}|\vec{A}|^{2}/E$ is preserved, as is the semi-major axis $a=\kappa/(2E)$ of the hyperbolic orbit. In the case $E=0$, $\vec{A}$ is preserved and $\vec{L}$ is shifted by a vector in the plane orthogonal to $\vec{A}$.
@@ 6. Concluding remarks
There are equivalent symmetry transformations that act in phase space, with the time $t$ being invariant. Their simplest formulation is given by considering the set of solutions $({\vec{r}}(t),{\vec{v}}(t))$ of the equations of motion (2.1). Under either of the dynamical symmetry transformations, each solution is mapped into a family of solutions
$({\vec{r}}(t),{\vec{v}}(t))\to({\vec{r}}^{\dagger}(t;\vec{\varepsilon}),{\vec{v}}^{\dagger}(t;\vec{\varepsilon}))$ (6.1)
with parameter $\vec{\varepsilon}$. To write the mapping in terms of the symmetry transformations, it is helpful to indicate the explicit dependence of these transformations (5.8) and (4.11) on the kinematic variables and the parameter: ${\vec{r}}^{*}={\vec{r}}^{*}[r,v;\varepsilon]$, ${\vec{v}}^{*}={\vec{v}}^{*}[r,v;\varepsilon]$, $t^{*}=t^{*}[t,r,v;\varepsilon]$. Now, for a fixed parameter $\vec{\epsilon}$, each solution represents a curve whose points $(t,{\vec{r}}(t),{\vec{v}}(t))$ get mapped to points with coordinates $\big(t^{*}[t,r(t),v(t);\varepsilon],{\vec{r}}^{*}[r(t),v(t);\varepsilon],{\vec{v}}^{*}[r(t),v(t);\varepsilon]\big)$ which lie on the curve $({\vec{r}}^{\dagger}(t;\vec{\varepsilon}),{\vec{v}}^{\dagger}(t;\vec{\varepsilon}))$, as shown in Fig. 6. This implies the relation
${\vec{r}}^{*}[r^{\dagger},v^{\dagger};-\varepsilon]={\vec{r}}(t^{*}(t,r^{\dagger},v^{\dagger};-\varepsilon)),\quad{\vec{v}}^{*}[r^{\dagger},v^{\dagger};-\varepsilon]={\vec{v}}(t^{*}(t,r^{\dagger},v^{\dagger};-\varepsilon))$ (6.2)
giving an implicit system of equations through which the coordinate values $({\vec{r}}^{\dagger},{\vec{v}}^{\dagger})$ can be determined at each time $t$. Hence, an implicit formulation of the mapping (6.1) is obtained.
[…]
All of the methods and results in the present work can be extended, firstly, to general central force dynamics. In particular, a generalization of the conserved LRL vector is known to exist for motion under any central force [13, 14, 15, 16, 17, 11] in $n>1$ dimensions. The physically relevant part of this vector is its direction in the plane of motion, which will give rise to an associated infinitesimal symmetry through Noether’s theorem and hence generate a corresponding transformation group of dynamical symmetries. A sequel paper [18] will study this group as well as the group containing the generalized LRL symmetry transformations.
@@
Appendix A Proof of Theorem 2
(1) The constants of motion $E$, $L^{i}$, $\Theta^{i}$ constitute 5 independent quantities, since they satisfy the two relations $L^{i}\Theta_{i}=0$ and $\Theta^{i}\Theta_{i}=1$. Together with $|r|$, they represent 6 degrees of freedom.
For any constant of motion $C$ in the Kepler problem, observe that
${\bf Y}_{\Theta^{j}}C=\tau^{j}D_{t}C+\hat{\bf X}_{\Theta^{j}}(C)=\hat{\bf X}_{\Theta^{j}}(C)$ (A.1)
holds on solutions of the equation of motion (2.1), since $D_{t}C=\mathcal{D}_{t}C=0$. This yields a differential equation
$\frac{d{C^{*}}}{d{\varepsilon_{j}}}=\hat{\bf X}_{\Theta^{j}}^{*}(C^{*})$ (A.2)
for the evolution of $C^{*}$ under the flow.
Use of expression (3.29) for the non-vanishing action of $\hat{\bf X}_{\Theta^{j}}$ on $E$, $L^{i}$, $\Theta^{i}$ yields a system of simple differential equations
$\frac{d{E^{*}}}{d{\varepsilon_{j}}}=0,\quad\frac{d{\Theta^{i}{}^{*}}}{d{\varepsilon_{j}}}=0,\quad\frac{d{L^{i}{}^{*}}}{d{\varepsilon_{j}}}=\epsilon^{ij}{}_{k}\Theta^{k}{}^{*}.$ (A.3)
Their solution is given by
$E^{*}=E,\quad\Theta^{i}{}^{*}=\Theta^{i},\quad L^{i}{}^{*}=\epsilon^{i}{}_{jk}\varepsilon^{j}\Theta^{k}+L^{i}.$ (A.4)
(2) Since $r^{i}$ and $v^{i}$ represent vectors that lie in the plane of motion and are linearly independent, they can be expressed as linear combinations of $\Theta^{i}$ and $\epsilon^{i}{}_{jk}L^{j}\Theta^{k}$ which are a pair of orthogonal vectors lying in the plane of motion. Their explicit form is given by
$\Theta^{i}$ $=\epsilon^{ijk}v_{j}L_{k}-\kappa|r|^{-1}r^{i}=(v^{j}v_{j}-\kappa|r|^{-1})r^{i}-(r^{j}v_{j})v^{i},$ (A.5)
$\epsilon^{i}{}_{jk}L^{j}\Theta^{k}$ $=v^{i}L^{j}L_{j}-\kappa|r|^{-1}\epsilon^{ijk}L_{j}r_{k}=(|r|^{2}(v^{j}v_{j}-\kappa|r|^{-1})-(r^{j}v_{j})^{2})v^{i}+(\kappa|r|^{-1}r^{j}v_{j})r^{i}.$ (A.6)
Thus,
$r^{i}=\alpha^{r}\Theta^{i}+\beta^{r}\epsilon^{i}{}_{kl}L^{k}\Theta^{l},\quad v^{i}=\alpha^{v}\Theta^{i}+\beta^{v}\epsilon^{i}{}_{kl}L^{k}\Theta^{l},$ (A.7)
where the coefficients in this basis expansion are given by
$\alpha^{r}=r_{j}\Theta^{j}=|L|^{2}-\kappa|r|,$ $\beta^{r}=|L|^{-2}r_{j}\epsilon^{j}{}_{kl}L^{k}\Theta^{l}=r^{j}v_{j},$ (A.8a)
$\alpha^{v}=v_{j}\Theta^{j}=-\kappa|r|^{-1}r^{j}v_{j},$ $\beta^{v}=|L|^{-2}v_{j}\epsilon^{j}{}_{kl}L^{k}\Theta^{l}=v^{j}v_{j}-\kappa|r|^{-1},$ (A.8b)
with
$|L|^{2}=L^{j}L_{j}=|r|^{2}v^{j}v_{j}-(r^{j}v_{j})^{2}.$ (A.9)
Now, via the expression (3.4) for energy, note that
$v^{j}v_{j}=2(E+\kappa|r|^{-1}),\quad r^{j}v_{j}=\pm\sqrt{|r|^{2}v^{j}v_{j}-|L|^{2}}=\pm\sqrt{2|r|(|r|E+\kappa)-|L|^{2}}$ (A.10)
where the $\pm$ denotes the sign of the radial velocity, ${\rm sgn}({\vec{r}}\cdot{\vec{v}})$. Substituting these expressions into the coefficients (A.8) then yields
$\alpha^{r}=|L|^{2}-\kappa|r|,\quad\beta^{r}={\rm sgn}({\vec{r}}\cdot{\vec{v}})|r|\sqrt{2(E+\kappa|r|^{-1})-(|r|^{-1}|L|)^{2}},$ (A.11a)
$\alpha^{v}=-{\rm sgn}({\vec{r}}\cdot{\vec{v}})\kappa\sqrt{2(E+\kappa|r|^{-1})-(|r|^{-1}|L|)^{2}},\quad\beta^{v}=2E+\kappa|r|^{-1},$ (A.11b)
which are expressed solely in terms of the constants of motion $E$ and $|L|$ and the invariant $|r|$.
Finally, under the flow (A.4), the expansion (A.7) yields the transformations (4.11b) on $r^{i}{}^{*}$ and $v^{i}{}^{*}$.
(3) The flow equation for $t^{*}$ now becomes
$\frac{d{t^{*}}}{d{\varepsilon_{j}}}=-\epsilon^{jik}r_{i}{}^{*}L_{k}{}^{*},$ (A.12)
which can be directly integrated as a quadrature. Substituting the transformation for for $r^{i}{}^{*}$ gives the transformation (4.11a).
Appendix B Proof of Theorem 3
(1) The 7 constants of motion $E$, $L^{i}$, $A^{i}$, together with $|r|$, represent 6 degrees of freedom, since $L^{i}A_{i}=0$ and $|A|^{2}=\kappa^{2}+2EL^{i}L_{i}$ hold identically. Their evolution under the flow (5.4) is given by the differential equation
$\frac{d{C^{*}}}{d{\varepsilon_{j}}}=\hat{\bf X}_{A^{j}}^{*}(C^{*}).$ (B.1)
Use of expressions (3.27)–(3.28) for the non-vanishing action of $\hat{\bf X}_{A^{j}}$ on $E$, $L^{i}$, and $A^{i}$ yields the system of 7 coupled differential equations
$\frac{d{E^{*}}}{d{\varepsilon_{j}}}=0,$ (B.2)
$\frac{d{L^{i}{}^{*}}}{d{\varepsilon_{j}}}=\epsilon^{ij}{}_{k}A^{k}{}^{*},\quad\frac{d{A^{i}{}^{*}}}{d{\varepsilon_{j}}}=-2E^{*}\epsilon^{ij}{}_{k}L^{k}{}^{*}.$ (B.3)
Observe that the solution of the first equation (B.2) is simply
$E^{*}=E.$ (B.4)
Then the remaining coupled equations (B.3) have a fairly simple form after the change of variable (3.12):
$\dfrac{d}{d\varepsilon^{j}}\begin{pmatrix}L^{i}{}^{*}\\ M^{i}{}^{*}\end{pmatrix}=\sqrt{2|E|}\epsilon^{ij}{}_{k}\begin{pmatrix}M^{k}{}^{*}\\ -{\rm sgn}(E)\,L^{k}{}^{*}\end{pmatrix}$ (B.5)
assuming $E\neq 0$.
In the case $E<0$, this system (B.5) is recognized to be a flow consisting of the twisted composition of an internal rotation with angular parameter $\sqrt{2|E|}$ and a spatial rotation through an angle
$|\varepsilon|=\varepsilon^{j}\varepsilon_{j},$ (B.6)
around an axis
$\hat{\varepsilon}^{j}=|\varepsilon|^{-1}\varepsilon^{j}.$ (B.7)
These two different rotations commute. The solution is given by
$\begin{pmatrix}L^{i}{}^{*}\\ M^{i}{}^{*}\end{pmatrix}=\exp\big(\sqrt{2|E|}\,\underline{J}\,\varepsilon^{j}\epsilon^{i}{}_{jk}\big)\begin{pmatrix}L^{k}\\ M^{k}\end{pmatrix},\quad\underline{J}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$. (B.8)
Expanding this exponential into a Taylor series and grouping the terms into even and odd powers of $\varepsilon^{j}$ yields the matrix operator
[…]
where
$\underline{R}^{i}{}_{k}(\phi,\hat{\varepsilon}^{j})=\cos(\phi)\ \underline{I}\,\delta^{i}{}_{k}+(1-\cos(\phi))\underline{I}\,\hat{\varepsilon}^{i}\hat{\varepsilon}_{k}+\sin(\phi)\underline{J}\,\hat{\varepsilon}_{j}\epsilon^{ji}{}_{k},\quad\underline{I}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$ (B.10)
represents a twisted rotation operator. Its parameters satisfy $\phi\in[0,2\pi)$, $\hat{\varepsilon}\in S^{2}$, where $S^{2}$ denotes the unit 2-sphere in $\mathbb{R}^{3}$. Hence,
$\begin{pmatrix}L^{i}{}^{*}\\ M^{i}{}^{*}\end{pmatrix}=\underline{R}^{i}{}_{k}(\sqrt{2|E|}\,|\varepsilon|,\hat{\varepsilon}^{j})\begin{pmatrix}L^{k}\\ M^{k}\end{pmatrix}$. (B.11)
In the case $E>0$, the flow (B.5) consists of the twisted composition of the preceding spatial rotation with angle $|\varepsilon|$ and axis $\hat{\varepsilon}^{j}$, and an internal hyperbolic rotation (boost) with parameter $\sqrt{2|E|}$. These two different rotations again commute. Thus, the solution is given by
$\begin{pmatrix}L^{i}{}^{*}\\ M^{i}{}^{*}\end{pmatrix}=\underline{\widetilde{R}}^{i}{}_{k}(\sqrt{2|E|}\,|\varepsilon|,\hat{\varepsilon}^{j})\begin{pmatrix}L^{k}\\ M^{k}\end{pmatrix}$ (B.12)
where
$\underline{\widetilde{R}}^{i}{}_{k}(\psi,\hat{\varepsilon}^{j})=\cosh(\psi)\ \underline{I}\,\delta^{i}{}_{k}+(1-\cosh(\psi))\underline{I}\,\hat{\varepsilon}^{i}\hat{\varepsilon}_{k}+\sinh(\psi)\underline{K}\,\hat{\varepsilon}_{j}\epsilon^{ji}{}_{k},\quad\underline{K}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$ (B.13)
represents a twisted boost-rotation operator. Its parameter satisfies $\psi\in(-\infty,\infty)$.
Last, in the case $E=0$, the equations (B.3) decouple, which directly yields the solution
$A^{i}{}^{*}=A^{i},\quad L^{i}{}^{*}=L^{i}+\epsilon^{i}{}_{jk}\varepsilon^{j}A^{k}.$ (B.14)
(2) The basis expansion (A.7)–(A.8) can be used again, where $\Theta^{i}$ is expressed in terms of $M^{i}$ in the case $E\neq 0$ and $A^{i}$ in the case $E=0$. Under the flow (B.4), (B.11), (B.12), (B.14), the resulting expansion then yields the transformations (5.8b) on $r^{i}{}^{*}$ and $v^{i}{}^{*}$.
(3) The flow equation for $t^{*}$ is determined entirely by $\tau^{j}$ which is the same as for the flow arising from the LRL direction vector. Recall that this differential equation can be directly integrated as a quadrature. Then substituting the transformation $r^{i}{}^{*}$ leads to the transformation (5.8a).
@@ Acknowledgements
S.C.A. is supported by an NSERC Discovery grant.
D. Holm and C. Stoica are thanked for helpful remarks.

### v3

@@ 5. LRL symmetry transformation group
The 6-dimensional symmetry transformation group acts as isometries in the space of constants of motion $(\vec{L},\vec{M})$, where $\vec{L}\cdot\vec{M}=0$ and $|\vec{L}|^{2}-{\rm sgn}(E)|\vec{M}|^{2}=\kappa^{2}/(2|E|)$, with $|\vec{L}|\neq 0$ and $|\vec{M}|\neq 0$. When $E<0$, the space is isomorphic to $S^{2}\times S^{1}\times S^{1}$, which has the symmetry group $SO(4)$. The space is isomorphic to $S^{2}\times S^{1}\times\mathbb{R}$ when $E>0$, and the symmetry group is then $SO(3,1)$.
In the case $E=0$, the LRL dynamical symmetry transformations (5.8) commute, forming an abelian Lie group isomorphic to $\mathbb{R}^{3}$, which also commutes with the $SO(3)$ group generated by the rotations.
@@
The physical meaning of the LRL dynamical symmetry transformations (5.8) is that they preserve energy, while angular momentum and the LRL vector are changed in a specific way that depends on the energy. In the case $E<0$, the vectors $\vec{L}$ and $\vec{M}=\sqrt{2|E|}^{-1}\vec{A}$ are rotated, preserving $|\vec{L}|^{2}+|\vec{M}|^{2}$ and $\vec{L}\cdot\vec{M}=0$, along with the period $T=2\pi/\sqrt{2|E|}$ and semi-major axis $a=\kappa/(2|E|)$ of the elliptical orbit. Likewise in the case $E>0$, $|\vec{L}|^{2}-\frac{1}{2}|\vec{A}|^{2}/E$ is preserved, as is the semi-major axis $a=\kappa/(2E)$ of the hyperbolic orbit. In the case $E=0$, $\vec{A}$ is preserved and $\vec{L}$ is shifted by a vector in the plane orthogonal to $\vec{A}$.
@@ 6. Concluding remarks
The methods and results in the present work can be extended, firstly, to general central force dynamics. In particular, a generalization of the conserved LRL vector is known to exist for any central force [12, 13, 14, 15, 16, 11] in $n>1$ dimensions. The physically relevant part of this vector is its direction in the plane of motion, which will give rise to an associated infinitesimal symmetry and thereby generate a transformation group of dynamical symmetries. A sequel paper [17] will study this group as well as the group containing the generalized LRL symmetry transformations and the temporal symmetry transformations associated to the additional integral of motion [11] which explicitly involves $t$.
@@ Acknowledgements
S.C.A. is supported by an NSERC Discovery grant. D. Holm and C. Stoica are thanked for helpful remarks.